OpenStudy (anonymous):

(4+radical10)(4-radical10)??? I Actually need steps On How to Do This, Test tomorrow, Please Help

OpenStudy (anonymous):

multiply each term in the first by each term in the second. This gives \[4*4-4*\sqrt{10}+4*\sqrt{10}-\sqrt{10}^2=16-10=6\]

OpenStudy (anonymous):

Do FOIL: (x-y)(x+y) = x*x +x*y - x*y - y*y = x^2 -y^2 If x = 4 & y = radical 10 The answer is 4^2 -(radical 10)^2 = 16-10 = 6

OpenStudy (anonymous):

What The Monkey Is FOIL?

OpenStudy (anonymous):

First items in the brackets Outside items in the brackets Inside items in the brackets Last items in the brackets....its a way to remember how to expand the equation

OpenStudy (anonymous):

The sqrt x+c * sqrt x-c ...that has a name....anyone recall what it is?

OpenStudy (anonymous):

What are you talking about: difference of squares?

OpenStudy (anonymous):

nope the square root version has its own name...

OpenStudy (anonymous):

nonsense. \[(a+b)(a-b)=a^2-b^2\] \[(4+\sqrt{10})(4-\sqrt{10})=4^2-\sqrt{10}^2=16-10=6\]

OpenStudy (anonymous):

'foil' is a trick math teachers play on students.

OpenStudy (anonymous):

conjugate...there it is sqrt x+c is the conjugate of sqrt x-c....probably the same for squares but it never came up. And I recall it coming up for complex numbers too now that I found it. i+c and i-c used to get rid of the complex number. =)